# VLM-E2E: Enhancing End-to-End Autonomous Driving with Multimodal Driver Attention Fusion

### v1

@@ Abstract
Human drivers adeptly navigate complex scenarios by utilizing rich attentional semantics, but the current autonomous systems struggle to replicate this ability, as they often lose critical semantic information when converting 2D observations into 3D space. In this sense, it hinders their effective deployment in dynamic and complex environments. Leveraging the superior scene understanding and reasoning abilities of Vision-Language Models (VLMs), we propose VLM-E2E, a novel framework that uses the VLMs to enhance training by providing attentional cues. Our method integrates textual representations into Bird’s-Eye-View (BEV) features for semantic supervision, which enables the model to learn richer feature representations that explicitly capture the driver’s attentional semantics. By focusing on attentional semantics, VLM-E2E better aligns with human-like driving behavior, which is critical for navigating dynamic and complex environments. Furthermore, we introduce a BEV-Text learnable weighted fusion strategy to address the issue of modality importance imbalance in fusing multimodal information. This approach dynamically balances the contributions of BEV and text features, ensuring that the complementary information from visual and textual modality is effectively utilized. By explicitly addressing the imbalance in multimodal fusion, our method facilitates a more holistic and robust representation of driving environments. We evaluate VLM-E2E on the nuScenes dataset and demonstrate its superiority over state-of-the-art approaches, showcasing significant improvements in performance.
Index Terms:
Bird’s eye view, multimodal information fusion, end-to-end autonomous driving, vision language models.
@@ I Introduction
Autonomous driving has witnessed remarkable progress in recent years [1, 2, 3], with significant advancements in key areas such as perception [4, 5, 6], motion prediction [7, 8, 9], and planning [10, 11]. These developments have laid a solid foundation for achieving more accurate and safer driving decisions. Among these, end-to-end (E2E) autonomous driving has emerged as a transformative paradigm, leveraging large-scale data to demonstrate impressive planning capabilities. By directly mapping raw sensor inputs to driving actions, E2E approaches bypass the need for handcrafted intermediate modules, enabling more flexible and scalable solutions. However, Despite these advancements, traditional end-to-end autonomous driving approaches predominantly predict future trajectories or control signals directly, without explicitly considering the driver’s attention to critical information such as traffic dynamics and navigation cues. E2E systems often struggle in complex and ambiguous scenarios due to their limited ability to reason about high-level semantics and contextual cues, such as traffic rules, driver attentions, and dynamic interactions. In contrast, human drivers rely on an attentional decision-making process, where attention to both the surrounding traffic environment and navigation guidance plays a critical role [12, 13, 14]. For instance, when approaching an intersection, human drivers naturally prioritize traffic signals, pedestrian movements, and lane markings, dynamically adjusting their focus based on the evolving scene.
@@
This limitation has spurred the integration of Vision-Language Models (VLMs) [15, 16, 17, 18] into autonomous driving frameworks. Trained on vast multimodal datasets, VLMs excel at tasks requiring high-level semantic reasoning, such as interpreting complex scenes, predicting dynamic interactions, and generating contextual descriptions. Their ability to leverage commonsense knowledge makes them particularly well-suited for addressing challenges in autonomous driving, such as understanding traffic rules, identifying vulnerable road users, and making safe decisions in ambiguous scenarios. By generating text-based descriptions of critical driving cues, VLMs can explicitly capture and prioritize regions of interest that align with human driver attention. This capability enables more human-like decision-making, particularly in safety-critical scenarios where attentional focus is paramount.
@@
How to integrate VLMs with E2E models? While most existing methods integrate VLMs with decision-making modules [19] or other high-level components [20, 21, 22], leveraging their semantic understanding capabilities to enhance decision processes, our approach introduces a novel integration strategy. Instead of limiting VLMs to decision modules, we combine them directly with the BEV module, which is widely used to represent and process spatial information from multiple perspectives in autonomous driving. By integrating VLMs into the BEV module, we enable BEV representations to incorporate both visual and textual features, resulting in richer and more semantic-aware spatial understanding. This integration allows the model to not only perceive geometric structures but also reason about high-level driver attentional semantics.
@@
How to fuse vision and text representations? Existing methods for fusing vision and text representations predominantly rely on attention-based mechanisms [23, 24], such as cross-attention or co-attention modules, to align and enhance interactions between modalities. While effective, these approaches often rely on predefined attention mechanisms that lack flexibility in adapting to varying task requirements and time and memory consuming. To address this issue, we propose a BEV-Text learnable weighted fusion strategy, where the importance of each modality is dynamically determined through a learnable weighting mechanism. This approach allows the model to adaptively emphasize visual or textual features based on their relevance to the task, leading to a more robust and context-aware multimodal representation. For instance, in scenarios requiring precise localization such as lane keeping, the model can prioritize BEV features, while in scenarios requiring high-level reasoning such as red lights, it can emphasize text features.
@@
How to represent driver attentional environment? To effectively model driver attentional environment, we propose a multimodal framework that leverages vision-language representations. First, we utilize front-view images captured from the driving scene as input to BLIP-2 [25] to generate initial textual descriptions of the environment. These descriptions provide a semantic understanding of key objects and events within the driver’s vision scope. To address the hallucination problem of VLMs, we further refine these textual representations using ground truth annotations and high-level maneuvering intentions. This refinement ensures that the generated text is not only accurate but also contextually aligned with the driving task. Finally, the refined text is encoded into a dense representation using a pre-trained CLIP [26] model, which aligns the textual information with visual features in a shared embedding space. This textual representation enables the model to capture driver attentional cues, such as focusing on pedestrians near crosswalks or traffic signals at intersections, leading to more human-like decision-making and better safety performance.
We evaluate VLM-E2E on the widely used nuScenes dataset, a comprehensive benchmark for autonomous driving research. Our experimental results demonstrate significant improvements over the baseline methods, highlighting the effectiveness of our approach in enhancing perception, decision-making, and overall driving performance. The key contributions of this work can be summarized as follows:
@@
  • To address the hallucination problem of VLMs, we incorporate semantic refinement of text descriptions generated from front-view images. By leveraging ground truth (GT) labels and high-level maneuvering intentions, we ensure that the textual representations are both accurate and highly relevant to the driving task, enhancing the model’s ability to reason about critical driving cues.
  • Extensive experiments on the nuScenes dataset demonstrate the superiority of VLM-E2E over existing methods. Our framework achieves significant improvements in handling complex driving scenarios, showcasing its ability to integrate geometric precision with high-level semantic reasoning for safer and more interpretable autonomous driving.
@@ II Related Work
II-A BEV Representation from Multi-view Cameras
BEV representation has emerged as a natural and ideal choice for planning and control tasks in autonomous driving [27, 28, 19, 29]. Unlike perspective views, BEV avoids issues such as occlusion and scale distortion while preserving the 3D spatial layout of the scene, making it highly suitable for tasks like path planning, object detection, and motion prediction. While LiDAR and HD maps can be easily represented in BEV, projecting vision inputs from camera views into BEV space remains a challenging problem due to the inherent complexity of perspective-to-BEV transformation [30].
Early approaches to BEV generation relied on geometric methods [31, 32], such as IPM [33], which assumes a flat ground plane to project 2D image pixels into BEV space. However, these methods struggle in complex environments where the ground plane assumption does not hold, such as uneven terrains or dynamic scenes. To address these limitations, learning-based methods have gained prominence. For instance, some works [34, 35] implicitly project image inputs into BEV using neural networks. However, the quality of these projections is often limited due to the lack of ground truth BEV data for supervision. Loukkal et al. [36] proposed an explicit projection method using homography between the image and BEV plane, but this approach is sensitive to calibration errors and environmental variations.
Recent advancements have introduced more sophisticated techniques for BEV generation. Methods like [37, 5] acquire BEV features through spatial cross-attention with pre-defined BEV queries, enabling end-to-end learning of perspective-to-BEV transformations. Notably, [4] and [38] have demonstrated impressive performance by leveraging estimated depth and camera intrinsics to perform the projection. These methods explicitly model the 3D geometry of the scene, resulting in more accurate and robust BEV representations.
II-B Multi-modal Information Fusion Mechanism
In recent years, attention-based fusion mechanisms and learnable fusion strategies have emerged as dominant paradigms for multi-modal information fusion, addressing the challenges of modality heterogeneity and imbalance. These approaches have demonstrated remarkable success in capturing cross-modal interactions and dynamically adapting to the relevance of each modality, making them particularly suitable for complex tasks such as autonomous driving and robotics.
Attention-based fusion mechanisms leverage the power of attention to model dependencies between modalities, enabling the model to focus on the most informative features. Transformer-based architectures [23, 24] have become a cornerstone of this approach, utilizing self-attention and cross-attention mechanisms to fuse features from different modalities. For instance, TransFuser [39] employs transformers to integrate visual and LiDAR features, achieving state-of-the-art performance in 3D object detection and scene understanding. Similarly, cross-modal attention networks [40] use attention to weigh the importance of visual and textual features, enhancing tasks such as image-text matching and visual question answering. These methods excel at capturing long-range dependencies and complex interactions between modalities. However, they often require significant computational resources, limiting their applicability in real-time systems.
On the other hand, learnable fusion mechanisms have gained traction for their ability to dynamically adjust the contribution of each modality based on task-specific requirements. These methods introduce learnable parameters, such as weights or coefficients, to adaptively fuse features during training. For example, Modality-Aware Fusion [41] proposes learnable coefficients to balance the importance of visual and LiDAR features, improving robustness in autonomous driving tasks. Another notable approach is Dynamic Fusion Networks [42], which use gating mechanisms to selectively combine modalities based on their relevance to the current context. These strategies are particularly effective in handling modality imbalance, where one modality may dominate due to its inherent information richness or task-specific importance. By dynamically adjusting the fusion process, learnable mechanisms ensure that all modalities contribute meaningfully to the final output, enhancing both performance and interpretability.
II-C End-to-end Autonomous Driving
End-to-end autonomous driving systems have demonstrated significant improvements in overall performance by jointly training all modules under a unified objective, thereby minimizing information loss across the pipeline. In recent years, unified frameworks such as ST-P3 [43] and UniAD [2] have pioneered vision-based E2E systems that seamlessly integrate perception, prediction, and planning modules, achieving state-of-the-art results in complex driving scenarios. Building on these advancements, subsequent research such as VAD [1] and VADv2 [44] introduced vectorized encoding methods to enhance the efficiency and scalability of scene representation, enabling more robust handling of dynamic environments.
More recently, methods such as Ego-MLP [45], BEV-Planner [46], and PARA-Drive [47] have explored novel design spaces within modular stacks, focusing on self-state modeling and innovative architectural designs to further enhance driving performance. These approaches have pushed the boundaries of E2E systems by incorporating richer representations of the ego vehicle’s state and its interactions with the environment.
In this work, we build upon ST-P3 by integrating driver attentional text information into the framework. By leveraging natural language descriptions of critical driving cues such as pedestrian crossing ahead or red traffic light, we enable the model to explicitly capture and prioritize regions of interest that align with human driver attention. This enhancement not only improves the interpretability of the system but also ensures that the model’s decisions are more closely aligned with human-like reasoning, particularly in safety-critical scenarios.
II-D Vision Language Models in Autonomous Driving
The integration of VLMs into autonomous driving systems has garnered significant attention due to their inherent capabilities in common sense knowledge, advanced reasoning, and interpretability. These attributes effectively address the limitations of traditional E2E models, making VLMs a promising avenue for enhancing driving systems. Recent research has explored various methodologies to harness VLMs for driving tasks, demonstrating substantial progress in this domain. For instance, Drive-with-LLMs [48] employs a Transformer network to encode ground-truth perception data into a latent space, which is subsequently processed by a Large Language Model (LLM) to predict future trajectories. Similarly, DriveGPT4 [49] leverages VLMs to interpret front-camera video inputs, generating planning control signals and providing natural language explanations for decision-making processes. Further advancements include DriveMLM [50], which validates the efficacy of VLM-based planning in closed-loop simulation environments [51], and ELM [52], which introduces large-scale pre-training of VLMs using cross-domain video data. These studies collectively underscore that the incorporation of diverse data sources and task-specific training significantly enhances VLM performance in driving-related tasks.
Moreover, several works have proposed specialized data collection strategies and datasets tailored for autonomous driving [53, 54, 55, 56], further accelerating the development and application of VLMs in this field. A notable contribution is DriveVLM [57], the first framework to seamlessly integrate VLMs with E2E models. In this approach, VLMs predict low-frequency trajectories, which are subsequently refined by the E2E model to generate the final planning trajectory. Additionally, Senna [20] generates high-level planning decisions in natural language, circumventing the need for VLMs to predict precise numerical values, thereby simplifying the decision-making process.
Despite these advancements, existing methods often overlook the design of driving-specific text prompts that capture the unique semantics of driving scenarios, particularly those incorporating driver attentional cues. These cues are crucial for achieving human-like decision-making in autonomous driving systems.
@@ III Methodology
In this section, we provide a detailed introduction to VLM-E2E, as illustrated in Fig. 2. The input scene information includes multi-view image sequences, GT, maneuvering, and user prompts. The front-view image, maneuvering, and user prompts are fed into the VLM-based Text Annotation Generation module to generate descriptive text annotations, while the multi-view images are processed by the visual encoding layer to produce BEV features. These text annotations are then passed to the Text Interaction Guidance Module, where they are encoded into text features using a pre-trained CLIP model. Subsequently, the BEV and text features are fused to support downstream tasks such as perception, prediction, and decision-making. In Section III-A, we introduce the design of VLM-based Text Annotation Generation in detail. Sections III-C and III-B focus on the design of the Text Interaction Guidance Module and Vision-based End-to-End Architecture, respectively.
@@ III-A1 Text Annotation
We encountered a challenge in determining the visual input. That is, selecting the right images from multiple cameras that can cover 360 degrees of the ego vehicle. Considering that we want to capture the driver’s attentional semantics when driving, the front view images usually contain the most relevant information required for most driving tasks. All-view images contain more distracting information that affects the system’s decision-making, so we choose to use only the front-view images to extract the attentional information. In addition, considering that the ego vehicle and its surroundings are in dynamic motion and the hallucination problem inherent in large models, we use the GT and maneuvering to refine the annotations of dynamic objects.
@@ III-B1 Text Encoder
Given a text input $T$ that provides semantic features to guide the BEV-Text fusion network toward achieving a specified fusion result, the text encoder and embedding within the text interaction guidance architecture are responsible for transforming this input into a text embedding. Among various VLMs, we adopt CLIP [26] due to its lightweight architecture and efficient text feature extraction capabilities. Compared to other VLMs, CLIP is computationally less demanding and produces text embeddings with a relatively small feature dimension of $77$, which significantly enhances the efficiency of subsequent BEV-Text feature fusion. We freeze the text encoder from CLIP to preserve its consistency and leverage its pretrained knowledge. This process can be formally expressed as:
@@
III-B2 BEV-Text Fusion
In the semantic interaction guidance module, semantic parameters interact through feature modulation and fusion features $s_{t}$, to obtain the effect of guidance. The feature modulation consists of scale scaling and bias control, which adjust the features from two perspectives, respectively. In particular, a residual connection is used to reduce the difficulty of network fitting, inspired by [58]. For simplicity, it can be described as:
@@ III-B2 BEV-Text Fusion
where $\odot$ denotes the Hadamard product. $x_{t}$ denotes the BEV feature of BEV-Text fusion, and $s_{t}$ denotes the BEV feature defined in Section III-C1.
@@ III-C1 Spatial Temporal BEV Perception
where $\hat{e}_{t}^{k}\in\mathbb{R}^{C\times D\times H_{e}\times W_{e}}$. Then, to transform the 2D perspective features into a 3D space, we employ a feature lifting module. This module uses camera intrinsic and extrinsic parameters to project the 2D features into a 3D voxel space. The 3D feature volume is then collapsed into a 2D BEV representation by aggregating features along the vertical axis to form the BEV view features $b_{t}\in\mathbb{R}^{C\times H\times W}$, with $(H,W)$ denotes the spatial size of BEV feature. This is achieved through attention-based aggregation, which preserves the most salient features while maintaining spatial consistency. The resulting BEV map provides a top-down view of the scene, encapsulating both geometric and semantic information.
@@
III-C2 Semantic Occupancy Prediction
The future prediction model is a convolutional gated recurrent unit network taking as input the current state $s_{t}$ and the latent variable $\eta_{t}$ sampled from the future distribution during training, or the present distribution $P$ for inference. It recursively predicts future states $(y_{t+1},\cdots,y_{t+l})$ with $l$ denotes the prediction horizon.
To model the inherent uncertainty in multi-modal future trajectories, we employ a conditional variational framework inspired by [60]. Present distribution $P(z|x_{t})$ is conditioned solely on the current state $x_{t}$. Future distribution $P_{f}(z|x_{t},y_{t+1:t+l})$ is augmented with ground-truth future observations $(y_{t+1},\cdots,y_{t+l})$. This distribution is parameterized as diagonal Gaussian with learnable mean $\mu\in\mathbb{R}^{M}$ and variance $\sigma^{2}\in\mathbb{R}^{M}$, where $M$ is the latent dimension.
$P(z|x_{t})=\mathcal{N}(\mu_{pres},\sigma^{2}_{press}),$ (8)
$P_{f}(z|x_{t},y_{t+1:t+l})=\mathcal{N}(\mu_{fut},\sigma^{2}_{fut})$ (9)
In the training phase, to ensure prediction consistency with observed futures while preserving multi-modal diversity, we sample $\eta_{t}$ from $P_{f}(z|x_{t},y_{t+1:t+l})$ and then optimize a mode-covering KL divergence loss.
$\mathcal{L}_{KL}=D_{KL}(P_{f}(z|x_{t},y_{t+1:t+F})||P(z|x_{t}))$ (10)
which encourages $P(z|x_{t})$ to encompass all plausible futures encoded in $P_{f}$. In the inference phase, future trajectories are generated by sampling from the present distribution $\eta_{t}\sim P(z|x_{t})$, where each sample $\eta_{t}$ represents a distinct future hypothesis.
This probabilistic formulation enables our model to generate diverse yet physically plausible futures while maintaining temporal consistency, crucial for handling ambiguous scenarios like unprotected left turns or pedestrian interactions.
The fusion features $x_{t}$ are processed by a multi-task decoder $D_{p}$ to generate instance-aware segmentation masks and motion predictions. The decoder outputs four key predictions: semantic segmentation, instance centerness, instance offset, and future instance flow, which collectively enable robust instance detection, segmentation, and tracking. The semantic segmentation head predicts pixel-wise semantic categories through a convolutional classifier. This provides a dense understanding of the scene layout and object categories. For instance segmentation, we adopt a hybrid center-offset formulation [61]. The instance centerness head outputs a heatmap $\mathcal{H}_{t}\in\mathbb{R}^{H\times W}$ indicating the likelihood of instance centers. During training, a Gaussian kernel is applied to suppress ambiguous regions and focus on high-confidence centers. The instance offset head predicts a vector field $\mathcal{O}_{t}\in\mathbb{R}^{2\times H\times W}$, where each vector points to its corresponding instance center. At inference, instance centers are extracted via non-maximum suppression (NMS) on $\mathcal{H}_{t}$. The future instance flow head predicts a displacement vector field $\mathcal{F}_{t}\in\mathbb{R}^{2\times H\times W}$ encoding the motion of dynamic agents over a future horizon $l$. This flow field is used to propagate instance centers across timesteps, ensuring temporal consistency. Specifically, detected instance centers $\{c_{i}^{t}\}$ are flow-warped to $t+1$ via $\hat{c}_{i}^{t+1}=c_{i}^{t}+\mathcal{F}_{t}(c_{i}^{t})$. The warped centers $\{\hat{c}_{i}^{t+1}\}$ are then matched to detected centers $c_{j}^{t+1}$ at $t+1$ using the Hungarian algorithm [62], which solves for optimal assignments based on pairwise IoU. This flow-based matching enables robust cross-frame association even under occlusions or abrupt motion changes.

### v2

@@ Abstract
Human drivers adeptly navigate complex scenarios by utilizing rich attentional semantics, but the current autonomous systems struggle to replicate this ability, as they often lose critical semantic information when converting 2D observations into 3D space. In this sense, it hinders their effective deployment in dynamic and complex environments. Leveraging the superior scene understanding and reasoning abilities of Vision-Language Models (VLMs), we propose VLM-E2E, a novel framework that uses the VLMs to enhance training by providing attentional cues. Our method integrates textual representations into Bird’s-Eye-View (BEV) features for semantic supervision, which enables the model to learn richer feature representations that explicitly capture the driver’s attentional semantics. By focusing on attentional semantics, VLM-E2E better aligns with human-like driving behavior, which is critical for navigating dynamic and complex environments. Furthermore, we introduce a BEV-Text learnable weighted fusion strategy to address the issue of modality importance imbalance in fusing multimodal information. This approach dynamically balances the contributions of BEV and text features, ensuring that the complementary information from visual and textual modalities is effectively utilized. By explicitly addressing the imbalance in multimodal fusion, our method facilitates a more holistic and robust representation of driving environments. We evaluate VLM-E2E on the nuScenes dataset and achieve significant improvements in perception, prediction, and planning over the baseline end-to-end model, showcasing the effectiveness of our attention-enhanced BEV representation in enabling more accurate and reliable autonomous driving tasks.
@@ I Introduction
Autonomous driving has witnessed remarkable progress in recent years [1, 2, 3], with significant advancements in key areas such as perception [4], motion prediction [5], and planning [6]. These developments have laid a solid foundation for achieving more accurate and safer driving decisions. Among these, end-to-end (E2E) autonomous driving has emerged as a transformative paradigm, leveraging large-scale data to demonstrate impressive planning capabilities. By directly mapping raw sensor inputs to driving actions, E2E approaches bypass the need for handcrafted intermediate modules, enabling more flexible and scalable solutions. However, Despite these advancements, traditional end-to-end autonomous driving approaches predominantly predict future trajectories or control signals directly, without explicitly considering the driver’s attention to critical information such as traffic dynamics and navigation cues. E2E systems often struggle in complex and ambiguous scenarios due to their limited ability to reason about high-level semantics and contextual cues, such as traffic rules, driver attention, and dynamic interactions. In contrast, human drivers rely on an attentional decision-making process, where attention to both the surrounding traffic environment and navigation guidance plays a critical role [7]. For instance, when approaching an intersection, human drivers naturally prioritize traffic signals, pedestrian movements, and lane markings, dynamically adjusting their focus based on the evolving scene.
@@
This limitation has spurred the integration of Vision-Language Models (VLMs) [8, 9] into autonomous driving frameworks. Trained on vast multimodal datasets, VLMs excel at tasks requiring high-level semantic reasoning, such as interpreting complex scenes, predicting dynamic interactions, and generating contextual descriptions. Their ability to leverage commonsense knowledge makes them particularly well-suited for addressing challenges in autonomous driving, such as understanding traffic rules, identifying vulnerable road users, and making safe decisions in ambiguous scenarios. By generating text-based descriptions of critical driving cues, VLMs can explicitly capture and prioritize regions of interest that align with human driver attention. This capability enables more human-like decision-making, particularly in safety-critical scenarios where attentional focus is paramount.
@@
How to integrate VLMs with E2E models? While most existing methods integrate VLMs with decision-making modules [10] or other high-level components [11], leveraging their semantic understanding capabilities to enhance decision processes, our approach introduces a novel integration strategy. Instead of limiting VLMs to decision modules, we combine them directly with the BEV module, which is widely used to represent and process spatial information from multiple perspectives in autonomous driving. By integrating VLMs into the BEV module, we enable BEV representations to incorporate both visual and textual features, resulting in richer and more semantic-aware spatial understanding. This integration allows the model to not only perceive geometric structures but also reason about high-level driver attentional semantics.
@@
How to fuse vision and text representations? Existing methods for fusing vision and text representations predominantly rely on attention-based mechanisms [12, 13], such as cross-attention or co-attention modules, to align and enhance interactions between modalities. While effective, these approaches often rely on predefined attention mechanisms that lack flexibility in adapting to varying task requirements and are time and memory-consuming. To address this issue, we propose a BEV-Text learnable weighted fusion strategy, where the importance of each modality is dynamically determined through a learnable weight mechanism. This approach allows the model to adaptively emphasize visual or textual features based on their relevance to the task, leading to a more robust and context-aware multimodal representation. For instance, in scenarios requiring precise localization such as lane keeping, the model can prioritize BEV features, while in scenarios requiring high-level reasoning, such as red lights, it can emphasize text features.
@@
How to represent driver attentional environment? To effectively model a driver-attentional environment, we propose a multimodal framework that leverages vision-language representations. First, we utilize front-view images captured from the driving scene as input to BLIP-2 [14] to generate initial textual descriptions of the environment. These descriptions provide a semantic understanding of key objects and events within the driver’s vision scope. To address the hallucination problem of VLMs, we further refine these textual representations using ground truth annotations and high-level maneuvering intentions. This refinement ensures that the generated text is not only accurate but also contextually aligned with the driving task. Finally, the refined text is encoded into a dense representation using a pre-trained CLIP [15] model, which aligns the textual information with visual features in a shared embedding space. This textual representation enables the model to capture driver attentional cues, such as focusing on pedestrians near crosswalks or traffic signals at intersections, leading to more human-like decision-making and better safety performance. The key contributions of this work can be summarized as follows:
@@
  • To address the hallucination problem of VLMs, we incorporate semantic refinement of textual annotations generated from front-view images. By leveraging ground truth (GT) labels and high-level maneuvering intentions, we ensure that the textual representations are both accurate and highly relevant to the driving task, enhancing the model’s ability to reason about critical driving cues.
  • Extensive experiments on the nuScenes dataset demonstrate the superiority of VLM-E2E over existing methods. Our framework achieves significant improvements in handling complex driving scenarios, showcasing its ability to integrate geometric precision with high-level semantic reasoning for safer and more reliable autonomous driving.
@@ II Related Work
II-A End-to-End Autonomous Driving
Recent advances in end-to-end autonomous driving systems have established vision-based frameworks like ST-P3 [16] and UniAD [2], which unify perception, prediction, and planning for improved scene understanding. Follow-up works such as VAD [1] and VADv2 [17] further enhanced dynamic environment handling through vectorized scene representations. Subsequent innovations like Ego-MLP [18] and BEV-Planner [19] introduce modular designs focusing on ego-state modeling and environment interactions.
II-B Vision Language Models in Autonomous Driving
VLMs have gained traction in autonomous driving for their ability to enhance reasoning and interpretability in end-to-end systems. Recent studies demonstrate diverse integration approaches. For instance, Drive-with-LLMs [20] processes perception data via transformers for trajectory prediction, while DriveGPT4 [21] generates control signals with natural language explanations. Frameworks like DriveMLM [22] and ELM [23] validate VLM-based planning through simulation and cross-domain pretraining.
Specialized datasets [24, 25] and hybrid systems such as DriveVLM [26], which refines VLM outputs into precise trajectories, and VLM-AD [10] and Senna [27], which simplify decision-making through text-based planning, highlight the field’s advancements. However, existing methods rarely address the design of driving-specific textual semantics, particularly those capturing driver-attentional cues essential for human-like situational awareness. Our work bridges this gap by integrating attentional textual prompts to prioritize safety-critical scenarios.
@@ III Methodology
In this section, we provide a detailed introduction to VLM-E2E, as illustrated in Fig. 2. The input scene information includes multi-view image sequences, GT, maneuvering, and user prompts. The front-view image, maneuvering, and user prompts are fed into the VLM-based Text Annotation Generation module to generate descriptive text annotations, while the multi-view images are processed by the visual encoding layer to produce BEV features. These text annotations are then passed to the Text Interaction Guidance Module, where they are encoded into text features using a pre-trained CLIP model. Subsequently, the BEV and text features are fused to support downstream tasks such as perception, prediction, and decision-making. In Section III-A, we introduce the design of VLM-based Text Annotation Generation in detail. Sections III-B and III-C focus on the design of the Text Interaction Guidance Module and Vision-based End-to-End Architecture, respectively.
@@ III-A1 Text Annotation
We encountered a challenge in determining the visual input. That is, selecting the right images from multiple cameras that can cover 360 degrees of the ego vehicle. Considering that we want to capture the driver’s attentional semantics when driving, the front view images usually contain the most relevant information required for most driving tasks. All-view images contain more distracting information that affects the system’s decision-making, so we choose to use only the front-view images to extract the attentional information. In addition, considering that the ego vehicle and its surroundings are in dynamic motion and the hallucination problem inherent in large models, we use the GT and maneuvering to refine the annotations of dynamic objects, inspired by [28].
@@ III-B1 Text Encoder
Given a text input $T$ that provides semantic features to guide the BEV-Text fusion network toward achieving a specified fusion result, the text encoder and embedding within the text interaction guidance architecture are responsible for transforming this input into a text embedding. Among various VLMs, we adopt CLIP [15] due to its lightweight architecture and efficient text feature extraction capabilities. Compared to other VLMs, CLIP is computationally less demanding and produces text embeddings with relatively small text tokens of $77$, which significantly enhances the efficiency of subsequent BEV-Text feature fusion. We freeze the text encoder from CLIP to preserve its consistency and leverage its pretrained knowledge. This process can be formally expressed as:
@@
As depicted in Fig. 2, in the text interaction guidance module, textual parameters interact through feature interaction and BEV features $s_{t}$, to obtain the effect of guidance. The feature modulation consists of scale scaling and bias control, which adjust the features from two perspectives, respectively. In particular, a residual connection is used to reduce the difficulty of network fitting, inspired by [29]. For simplicity, it can be described as:
@@
where $\odot$ denotes the Hadamard product. $x_{t}$ denotes the fused BEV feature, and $s_{t}$ denotes the BEV feature defined in Section III-C1.
@@ III-C1 Spatial Temporal BEV Perception
where $\hat{e}_{t}^{k}\in\mathbb{R}^{C\times D\times H_{e}\times W_{e}}$. Then, to transform the 2D perspective features into a 3D space, we employ a feature lifting module. This module uses camera intrinsic and extrinsic parameters to project the 2D features into a 3D voxel space. The 3D feature volume is then collapsed into a 2D BEV representation by aggregating features along the vertical axis to form the BEV view features $b_{t}\in\mathbb{R}^{C\times H\times W}$, with $(H,W)$ denoting the spatial size of BEV feature. This is achieved through attention-based aggregation, which preserves the most salient features while maintaining spatial consistency. The resulting BEV map provides a top-down view of the scene, encapsulating both geometric and semantic information.